OpenStudy (anonymous):

Given f(x) = 3x - 1 and g(x)= -x + 6, find f(-2) + g(5). -6, 6, or 8?

OpenStudy (anonymous):

you're gonna plug in -2 into the f(x) and then you're gonna plug in 5 into the g(x) and then the two answers that you get you'll add and it should be one of the choices.

OpenStudy (anonymous):

And?

OpenStudy (usukidoll):

O_O plug in f(-2) in the f function plug in g(5) in the g function and add the results up

OpenStudy (anonymous):

f(x)= 3(-2) -1= -6-1= -7 g(x)= -5+6=1 f(-2)+g(5)= -7+1=-6

OpenStudy (anonymous):

Then you get your answer?

OpenStudy (anonymous):

Thanks Dwade!

OpenStudy (anonymous):

no problem :)